the base of an isosceles triangle and one of its legs have lengths that are consecutive integers. the leg is longer than the base. the perimeter of the triangle is 20. find the length of each side of the triangle.
Answer by checkley71(8403)

LET X BE THE SIDES OF THE ISOSCELES TRIANGLE.
THE BASE IS (X-1).
PERIMETER=2*SIDES+BASE
20=2X+(X-1)
20=3X-1
3X=20+1
3X=21
X=7 FOR THE LENGTH OF THE SIDES.
THUS THE BASE=7-1=6
PROOF:
20=2*7+(7-1)
20=14+6
20=20